OpenStudy (anonymous):

simplify the rational expression. state any excluded values 2 x/-5x+x^2

OpenStudy (dtan5457):

Hint: Factor -5x+x^2

OpenStudy (dtan5457):

Then it becomes obvious lol

OpenStudy (dtan5457):

\[\frac{ 2x }{ -5x+x^2 }=\frac{ 2x }{ x(x-5) }=?\] If you have two of the same values on the numerator and denominator, you can cross them out.